# Partial identification of principal causal effects under violations of principal ignorability

### v2

@@ 7 Extensions to Bayesian nonparametric modeling
$\begin{split}Y_{i}\mid\boldsymbol{U}_{i},\boldsymbol{X_{i}},T_{i}=t&\sim\mu_{yt}(\boldsymbol{X}_{i},S_{i})+\mu_{yct}(\boldsymbol{X}_{i},\boldsymbol{U}_{i})+\epsilon_{i},\\ \boldsymbol{U}_{i}\mid\boldsymbol{X}_{i}&\sim N(\boldsymbol{\mu}_{s}(\boldsymbol{X}),\boldsymbol{\Sigma}_{s}),\end{split}$ (6)
@@ 7.1 Partial identification related to extended sign issues
Theorem 3 shows that $\mu_{yct}(\boldsymbol{X},\boldsymbol{U})$ is unidentifiable, and therefore $E\{Y(t)\mid S(t),S(1-t),\boldsymbol{X}\}$ and any PCEs of interest are also unidentifiable. Because of the unobserved nature of $S(1-t)$, the flexibility of nonparameteric outcome models allows various possible ways for $S(1-t)$ to contribute to the observed outcome. These plausible outcome models can lead to completely different or even contradictory PCEs. For instance, if $(S(t),S(1-t))$ follows a bivariate Gaussian distribution, then one such invariant transformation can be constructed as $M:s\to 2m_{1-t}(s_{t},\boldsymbol{x})-s$, where $m_{1-t}(s_{t},\boldsymbol{x})$ denotes $E\{S(1-t)\mid S(t)=s_{t},\boldsymbol{X}=\boldsymbol{x}\}$. Clearly, $M$ is invertible and $M^{-1}=M$. Partial identification arises if $\mu_{yct}\{\boldsymbol{x},s_{t},2m_{1-t}(s_{t},\boldsymbol{x})-s_{1-t}\}\neq\mu_{yct}(\boldsymbol{x},s_{t},s_{1-t})$, which is typically the case. To provide intuition for this, this inequality would hold in the linear model in (1) if $\beta_{t,1-t}$ is not equal to 0, or equivalently, when principal ignorability fails.
Although we assume a bivariate Gaussian distribution for $(S(t),S(1-t))$ in Model (6), Theorem 3 does not rely on this assumption and holds for $(S(t),S(1-t))$ under any distribution. Theorem 3 generalizes the sign issues leading to partial identification in the linear model setting to a more general partial identification result based on a class of invariant transformations. Usually, the number of transformations in this class is infinite. However, most are artificial, resulting in complex and unusual potential outcome models $Y^{\prime}(t)$, which can easily be ruled out under reasonable modeling constraints. For cases like the sign issue, one can use domain knowledge to determine the correct transformation or employ reasonably uninformative priors to rule out other, incorrect outcome models.
@@ 8 Discussion
.1 Proofs Relating to Model (1)
@@
where $\mu^{\prime}_{yt}=\boldsymbol{\beta}_{t}^{T}\boldsymbol{\phi}_{i}+\lambda_{t}$, $\zeta_{t}=\sigma_{y}^{2}+\boldsymbol{\beta}_{t}^{T}\boldsymbol{\Sigma}_{s}\boldsymbol{\beta}_{t}$, and $\psi_{t}=\sigma_{st}\beta_{tt}+\rho\sigma_{s,1-t}\beta_{t,1-t}$.
@@
$p(\cdot)=\prod_{i=1}^{n}\bigg{(}(2\pi\sigma_{s,t_{i}})^{-1}(\zeta_{t_{i}}-\psi_{t_{i}}^{2})^{-1/2}\exp\bigg{(}-\frac{1}{2}\bigg{(}\frac{(y_{i}^{\prime\prime}-\psi_{t_{i}}s_{i}^{\prime\prime}/\sigma_{s,t_{i}})^{2}}{\zeta_{t_{i}}-\psi_{t_{i}}^{2}}+\frac{s_{i}^{\prime\prime 2}}{\sigma_{s,t_{i}}^{2}}\bigg{)}\bigg{)}\bigg{)},$
@@
$q_{n}(\rho)=-\frac{1}{2}\sum\limits_{i=1}^{n}\left(\log(\zeta_{t_{i}}-\psi_{t_{i}}^{2})+\frac{(y_{i}^{\prime\prime}-\psi_{t_{i}}s_{i}^{\prime\prime}/\sigma_{s,t_{i}})^{2}}{\zeta_{t_{i}}-\psi_{t_{i}}^{2}}+s_{i}^{\prime\prime 2}/\sigma_{s,t_{i}}^{2}\right)+\log(f(\rho)),$
@@
Then, $q_{n}(\rho)=\sum\limits_{i=1}^{n}(t_{i}l_{i}(\rho\mid 1)+(1-t_{i})l_{i}(\rho\mid 0))+\log(f(\rho))/n$. Let $n_{0}=\sum\limits_{i=1}^{n}(1-t_{i})$ and $n_{1}=\sum\limits_{i=1}^{n}t_{i}$. By the positivity assumption, $n_{0}\to\infty$ and $n_{1}\to\infty$, as $n\to\infty$. Within the control group ($T_{i}=0$) or the treatment group ($T_{i}=1$), $(Y^{\prime}_{i},S^{\prime}_{i})$ are i.i.d.
Let $\hat{\rho}_{0}=\operatorname*{arg\,max}_{\rho}(\sum\limits_{i=1}^{n}(1-t_{i})l_{i}(\rho\mid 0)/n_{0})$ and $\hat{\rho}_{1}=\operatorname*{arg\,max}_{\rho}(\sum\limits_{i=1}^{n}t_{i}l_{i}(\rho\mid 1)/n_{1})$. Next, we show that $\hat{\rho}_{0}\to{{\rho}^{\ast}}$ and $\hat{\rho}_{1}\to{{\rho}^{\ast}}$ in probability as $n\to\infty$. By Proposition 7.1 in [14], we need to show that they satisfy the identification condition and the uniform convergence condition within both groups.
@@
    As long as $\beta_{01}\neq 0$, we have $02\sigma_{s1}^{2}\beta_{01}^{2}/(\zeta_{0}-\psi_{0}^{2})(-\sigma_{s1}^{2}\beta_{01}^{2}\rho^{2}/(\zeta_{0}-\psi_{0}^{2})-1)<$ 0, where $\zeta_{0}-\psi_{0}^{2}=\sigma_{y}^{2}+(1-\rho^{2})\beta_{01}^{2}\sigma_{s1}^{2}>0$. Thus, $E(l_{i}(\rho\mid 0))$ is strictly increasing on $[-1,{{\rho}^{\ast}})$ and strictly decreasing on $({{\rho}^{\ast}},1]$, indicating that $E(l_{i}(\rho\mid 0))$ is uniquely maximized on $[-1,1]$ at ${{\rho}^{\ast}}$.
@@
    Note that $\text{Var}\left(\sum\limits_{i=1}^{n}(1-t_{i})l_{i}(\rho\mid 0)/n_{0}\right)=\text{Var}(l_{i}(\rho\mid 0))/n_{0}=4/n_{0}$. By Chebyshev’s inequality, uniform convergence holds.
@@
Since $q_{n}({\rho})=\sum\limits_{i=1}^{n}\left(t_{i}l_{i}(\rho\mid 1)/n_{0}(n_{0}/n)+(1-t_{i})l_{i}(\rho\mid 0)/n_{1}(n_{1}/n)\right)+\log(f(\rho))/n$, the above provides intuition that $\hat{\rho}_{B}$ (posterior mode) converges to ${{\rho}^{\ast}}$. To complete our proof rigorously, we will show that for any $\epsilon>0$, $P(|\hat{\rho}_{B}-{{\rho}^{\ast}}|>\epsilon)\to 0$ as $n\to\infty$. If $f(\rho)$ is continuous or bounded, then $\log(f(\rho))/n\to 0$ uniformly as $n\to\infty$. Since $\sum\limits_{t_{i}=t}l_{i}(\rho\mid t)/n_{t}$ uniformly converges to $E(l_{i}(\rho\mid t))$, we have $|q_{n}({\rho})-E(l_{i}(\rho\mid 0))(n_{0}/n)-E(l_{i}(\rho\mid 1))(n_{1}/n)|\to 0$ uniformly in probability, and $P(q_{n}(\rho)\geq q_{n}({{\rho}^{\ast}}):|\rho-{{\rho}^{\ast}}|>\epsilon)\to 0$, that is, $P(|\hat{\rho}_{B}-{{\rho}^{\ast}}|>\epsilon)\to 0$.
.1.2 Proof of Theorem 2
@@
$\begin{split}E(\rho\mid\mathcal{D})&\approx\frac{\int_{\rho^{*}-\delta_{1}}^{\rho^{*}+\delta_{2}}\rho\exp(q_{n}(\hat{\rho}_{B})+\frac{1}{2}q_{n}^{\prime\prime}(\hat{\rho}_{B})(\rho-\hat{\rho}_{B})^{2}+o((\rho-\hat{\rho}_{B})^{2})d\rho}{\int_{\rho^{*}-\delta_{1}}^{\rho^{*}+\delta_{2}}\exp(q_{n}(\hat{\rho}_{B})+\frac{1}{2}q_{n}^{\prime\prime}(\hat{\rho}_{B})(\rho-\hat{\rho}_{B})^{2}+o((\rho-\hat{\rho}_{B})^{2})d\rho}\\ &=\hat{\rho}_{B},\end{split}$
@@
$\begin{split}E(\rho^{2}\mid\mathcal{D})&\approx\frac{\int_{\rho^{*}-\delta_{1}}^{\rho^{*}+\delta_{2}}\rho^{2}\exp(q_{n}(\hat{\rho}_{B})+\frac{1}{2}q_{n}^{\prime\prime}(\hat{\rho}_{B})(\rho-\hat{\rho}_{B})^{2}+o((\rho-\hat{\rho}_{B})^{2})d\rho}{\int_{\rho^{*}-\delta_{1}}^{\rho^{*}+\delta_{2}}\exp(q_{n}(\hat{\rho}_{B})+\frac{1}{2}q_{n}^{\prime\prime}(\hat{\rho}_{B})(\rho-\hat{\rho}_{B})^{2}+o((\rho-\hat{\rho}_{B})^{2})d\rho}\\ &=\hat{\rho}_{B}^{2}-\frac{1}{q_{n}^{\prime\prime}(\hat{\rho}_{B})}.\end{split}$
Therefore, an approximation of the posterior variance is $-\frac{1}{q_{n}^{\prime\prime}(\hat{\rho}_{B})}$. By the Strong Law of Large Numbers (SLLN), we have
$\sum\limits_{t_{i}=t}l^{\prime\prime}_{i}(\rho\mid t)n_{t}^{-1}\to E(l^{\prime\prime}_{i}(\rho\mid t)),\text{ almost surely,}$,
@@
Since $q^{\prime\prime}_{n}(\rho)/n=\sum\limits_{t_{i}=0}(l^{\prime\prime}_{i}(\rho\mid 0)/n_{0})(n_{0}/n)+\sum\limits_{t_{i}=1}(l^{\prime\prime}_{i}(\rho\mid 1)/n_{1})(n_{1}/n)$, then an asymptomatic approximation of $q_{n}^{\prime\prime}(\rho)/n$ can be
@@
where $\bar{T}=\sum\limits_{i=1}^{n}T_{i}/n$.
@@
$\begin{cases}&\mu_{y0}^{\prime}=\lambda_{0}+\beta_{00}\phi_{0}+\beta_{01}\phi_{1}\\ &\mu_{y1}^{\prime}=\lambda_{1}+\beta_{10}\phi_{0}+\beta_{11}\phi_{1}\\ &\zeta_{0}=\sigma_{y}^{2}+\boldsymbol{\beta}_{0}^{T}\boldsymbol{\Sigma}_{s}\boldsymbol{\beta}_{0}\\ &\zeta_{1}=\sigma_{y}^{2}+\boldsymbol{\beta}_{1}^{T}\boldsymbol{\Sigma}_{s}\boldsymbol{\beta}_{1}\\ &\psi_{0}=\sigma_{s0}\beta_{00}+\rho\sigma_{s1}\beta_{01}\\ &\psi_{1}=\sigma_{s1}\beta_{11}+\rho\sigma_{s0}\beta_{10}\end{cases}$ (10)
@@
$\begin{split}\frac{(\text{Cov}(Y(t),S(1-t)\mid S(t),\boldsymbol{X}))^{2}}{\text{Var}(S(1-t)\mid S(t),\boldsymbol{X})\text{Var}(Y\mid S(t),\boldsymbol{X})}&\leq\frac{(\text{Cov}(Y(t),S(t)\mid\boldsymbol{X}))^{2}}{\text{Var}(Y\mid\boldsymbol{X})\text{Var}(S(t)\mid\boldsymbol{X})}\\ \frac{\beta_{t,1-t}^{2}\sigma^{4}_{s,1-t}(1-\rho^{2})^{2}}{\sigma^{2}_{s,1-t}(1-\rho^{2})(\sigma_{y}^{2}+\beta_{t,1-t}^{2}\sigma^{2}_{s,1-t}(1-\rho^{2}))}&\leq\frac{(\beta_{tt}\sigma_{st}+\rho\beta_{t,1-t}\sigma_{s,1-t})^{2}}{\sigma_{st}^{2}(\sigma_{y}^{2}+\boldsymbol{\beta}_{t}^{T}\boldsymbol{\Sigma}_{s}\boldsymbol{\beta}_{t})}\\ \end{split}$
@@
.2 Proofs Relating to Model (6)
.2.1 Proof of Theorem 3
@@
Although $M$ could relate to $S(t)$ and $\boldsymbol{X}$, we do not include them in the $M$ for simplicity of notation. Let $S^{\prime}(1-t)$ denote $M(S(1-t))$ and $p_{\epsilon}$ denote the pdf of $\epsilon$. Throughout, for simplicity of notation we let $p\big{(}Y^{\prime}(t)=y)$ denote the pdf of $Y^{\prime}(t)$ evaluated at $y$, with similar notation for $S(t)$. The the conditional distribution of $Y^{\prime}(t)$ given $S(t)$ and $\boldsymbol{X}$ is
[…]
@@
.2.2 Proof of an equivalent expression for Assumption 8

### v3

@@ 7 Extensions to Bayesian nonparametric modeling
[…]
@@ 7.1 Partial identification related to extended sign issues
Theorem 3 shows that $\mu_{yct}(\boldsymbol{X},\boldsymbol{U})$ is unidentifiable, and therefore $E\{Y(t)\mid S(t),S(1-t),\boldsymbol{X}\}$ and any PCEs of interest are also unidentifiable. Because of the unobserved nature of S(1t)S(1-t), the flexibility of nonparameteric outcome models allows various possible ways for S(1t)S(1-t) to contribute to the observed outcome. These plausible outcome models can lead to completely different or even contradictory PCEs. For instance, if (S(t),S(1t))(S(t),S(1-t)) follows a bivariate Gaussian distribution, then one such invariant transformation can be constructed as $M:s\to 2m_{1-t}(s_{t},\boldsymbol{x})-s$, where $m_{1-t}(s_{t},\boldsymbol{x})$ denotes $E\{S(1-t)\mid S(t)=s_{t},\boldsymbol{X}=\boldsymbol{x}\}$. Clearly, $M$ is invertible and $M^{-1}=M$. Partial identification arises if $\mu_{yct}\{\boldsymbol{x},s_{t},2m_{1-t}(s_{t},\boldsymbol{x})-s_{1-t}\}\neq\mu_{yct}(\boldsymbol{x},s_{t},s_{1-t})$, which is typically the case. To provide intuition for this, this inequality would hold in the linear model in (1) if $\beta_{t,1-t}$ is not equal to 0, or equivalently, when principal ignorability fails.
Although we assume a bivariate Gaussian distribution for (S(t),S(1t))(S(t),S(1-t)) in Model (5), Theorem 3 does not rely on this assumption and holds for (S(t),S(1t))(S(t),S(1-t)) under any distribution. Theorem 3 generalizes the sign issues leading to partial identification in the linear model setting to a more general partial identification result based on a class of invariant transformations. Usually, the number of transformations in this class is infinite. However, most are artificial, resulting in complex and unusual potential outcome models $Y^{\prime}(t)$, which can easily be ruled out under reasonable modeling constraints. For cases like the sign issue, one can use domain knowledge to determine the correct transformation or employ reasonably uninformative priors to rule out other, incorrect outcome models.
@@ 8 Discussion
Appendix A Proofs
A.1 Proofs Relating to Model (1)
@@ A.1.1 Proof of Theorem 1
[…]
@@
$p(\cdot)=\prodop\displaylimits_{i=1}^{n}\bigg((2\pi\sigma_{s,t_{i}})^{-1}(\zeta_{t_{i}}-\psi_{t_{i}}^{2})^{-1/2}\exp\bigg(-\frac{1}{2}\bigg(\frac{(y_{i}^{\prime\prime}-\psi_{t_{i}}s_{i}^{\prime\prime}/\sigma_{s,t_{i}})^{2}}{\zeta_{t_{i}}-\psi_{t_{i}}^{2}}+\frac{s_{i}^{\prime\prime 2}}{\sigma_{s,t_{i}}^{2}}\bigg)\bigg)\bigg),$
@@
$q_{n}(\rho)=-\frac{1}{2}\sumop\displaylimits\limits_{i=1}^{n}\left(\log(\zeta_{t_{i}}-\psi_{t_{i}}^{2})+\frac{(y_{i}^{\prime\prime}-\psi_{t_{i}}s_{i}^{\prime\prime}/\sigma_{s,t_{i}})^{2}}{\zeta_{t_{i}}-\psi_{t_{i}}^{2}}+s_{i}^{\prime\prime 2}/\sigma_{s,t_{i}}^{2}\right)+\log(f(\rho)),$
@@
Then, $q_{n}(\rho)=\sumop\displaylimits\limits_{i=1}^{n}(t_{i}l_{i}(\rho\mid 1)+(1-t_{i})l_{i}(\rho\mid 0))+\log(f(\rho))/n$. Let $n_{0}=\sumop\displaylimits\limits_{i=1}^{n}(1-t_{i})$ and $n_{1}=\sumop\displaylimits\limits_{i=1}^{n}t_{i}$. By the positivity assumption, $n_{0}\to\infty$ and $n_{1}\to\infty$, as $n\to\infty$. Within the control group ($T_{i}=0$) or the treatment group ($T_{i}=1$), $(Y^{\prime}_{i},S^{\prime}_{i})$ are i.i.d.
Let $\hat{\rho}_{0}=\operatorname*{arg\,max}_{\rho}(\sumop\displaylimits\limits_{i=1}^{n}(1-t_{i})l_{i}(\rho\mid 0)/n_{0})$ and $\hat{\rho}_{1}=\operatorname*{arg\,max}_{\rho}(\sumop\displaylimits\limits_{i=1}^{n}t_{i}l_{i}(\rho\mid 1)/n_{1})$. Next, we show that $\hat{\rho}_{0}\to{{\rho}^{\ast}}$ and $\hat{\rho}_{1}\to{{\rho}^{\ast}}$ in probability as $n\to\infty$. By Proposition 7.1 in [16], we need to show that they satisfy the identification condition and the uniform convergence condition within both groups.
@@
    As long as $\beta_{01}\neq 0$, we have $2\sigma_{s1}^{2}\beta_{01}^{2}/(\zeta_{0}-\psi_{0}^{2})(-\sigma_{s1}^{2}\beta_{01}^{2}\rho^{2}/(\zeta_{0}-\psi_{0}^{2})-1)<0$, where $\zeta_{0}-\psi_{0}^{2}=\sigma_{y}^{2}+(1-\rho^{2})\beta_{01}^{2}\sigma_{s1}^{2}>0$. Thus, $E(l_{i}(\rho\mid 0))$ is strictly increasing on $1,{{\rho}^{\ast}})$ and strictly decreasing on $({{\rho}^{\ast}},1]$, indicating that $E(l_{i}(\rho\mid 0))$ is uniquely maximized on [1,1][-1,1] at $\rho$$^{\ast}$}.
@@
    Note that $\text{Var}\left(\sumop\displaylimits\limits_{i=1}^{n}(1-t_{i})l_{i}(\rho\mid 0)/n_{0}\right)=\text{Var}(l_{i}(\rho\mid 0))/n_{0}=4/n_{0}$. By Chebyshev’s inequality, uniform convergence holds.
@@
Since $q_{n}({\rho})=\sumop\displaylimits\limits_{i=1}^{n}\left(t_{i}l_{i}(\rho\mid 1)/n_{0}(n_{0}/n)+(1-t_{i})l_{i}(\rho\mid 0)/n_{1}(n_{1}/n)\right)+\log(f(\rho))/n$, the above provides intuition that $\hat{\rho}_{B}$ (posterior mode) converges to $\rho$$^{\ast}$}. To complete our proof rigorously, we will show that for any $\epsilon>0$, $P(|\hat{\rho}_{B}-{{\rho}^{\ast}}|>\epsilon)\to 0$ as $n\to\infty$. If $f(\rho)$ is continuous or bounded, then $\log(f(\rho))/n\to 0$ uniformly as $n\to\infty$. Since $\sumop\displaylimits\limits_{t_{i}=t}l_{i}(\rho\mid t)/n_{t}$ uniformly converges to $E(l_{i}(\rho\mid t))$, we have $|q_{n}({\rho})-E(l_{i}(\rho\mid 0))(n_{0}/n)-E(l_{i}(\rho\mid 1))(n_{1}/n)|\to 0$ uniformly in probability, and $P(q_{n}(\rho)\geq q_{n}({{\rho}^{\ast}}):|\rho-{{\rho}^{\ast}}|>\epsilon)\to 0$, that is, $P(|\hat{\rho}_{B}-{{\rho}^{\ast}}|>\epsilon)\to 0$.
A.1.2 Proof of Theorem 2
@@
$\begin{split}E(\rho\mid\mathcal{D})&\approx\frac{\intop\nolimits_{\rho^{*}-\delta_{1}}^{\rho^{*}+\delta_{2}}\rho\exp(q_{n}(\hat{\rho}_{B})+\frac{1}{2}q_{n}^{\prime\prime}(\hat{\rho}_{B})(\rho-\hat{\rho}_{B})^{2}+o((\rho-\hat{\rho}_{B})^{2})d\rho}{\intop\nolimits_{\rho^{*}-\delta_{1}}^{\rho^{*}+\delta_{2}}\exp(q_{n}(\hat{\rho}_{B})+\frac{1}{2}q_{n}^{\prime\prime}(\hat{\rho}_{B})(\rho-\hat{\rho}_{B})^{2}+o((\rho-\hat{\rho}_{B})^{2})d\rho}\\ &=\hat{\rho}_{B},\end{split}$
@@
$\begin{split}E(\rho^{2}\mid\mathcal{D})&\approx\frac{\intop\nolimits_{\rho^{*}-\delta_{1}}^{\rho^{*}+\delta_{2}}\rho^{2}\exp(q_{n}(\hat{\rho}_{B})+\frac{1}{2}q_{n}^{\prime\prime}(\hat{\rho}_{B})(\rho-\hat{\rho}_{B})^{2}+o((\rho-\hat{\rho}_{B})^{2})d\rho}{\intop\nolimits_{\rho^{*}-\delta_{1}}^{\rho^{*}+\delta_{2}}\exp(q_{n}(\hat{\rho}_{B})+\frac{1}{2}q_{n}^{\prime\prime}(\hat{\rho}_{B})(\rho-\hat{\rho}_{B})^{2}+o((\rho-\hat{\rho}_{B})^{2})d\rho}\\ &=\hat{\rho}_{B}^{2}-\frac{1}{q_{n}^{\prime\prime}(\hat{\rho}_{B})}.\end{split}$
Therefore, an approximation of the posterior variance is $\frac{1}{q_{n}^{\prime\prime}(\hat{\rho}_{B})}$. By the Strong Law of Large Numbers (SLLN), we have
$\sumop\displaylimits\limits_{t_{i}=t}l^{\prime\prime}_{i}(\rho\mid t)n_{t}^{-1}\to E(l^{\prime\prime}_{i}(\rho\mid t)),\text{ almost surely,}$
@@
Since $q^{\prime\prime}_{n}(\rho)/n=\sumop\displaylimits\limits_{t_{i}=0}(l^{\prime\prime}_{i}(\rho\mid 0)/n_{0})(n_{0}/n)+\sumop\displaylimits\limits_{t_{i}=1}(l^{\prime\prime}_{i}(\rho\mid 1)/n_{1})(n_{1}/n)$, then an asymptomatic approximation of $q_{n}^{\prime\prime}(\rho)/n$ can be
@@
where $\bar{T}=\sumop\displaylimits\limits_{i=1}^{n}T_{i}/n$.
@@ A.1.3 Proof of Proposition 1
[…]
@@ A.1.6 Proof of Proposition 3
[…]
@@
A.2 Proofs Relating to Model (5)
A.2.1 Proof of Theorem 3
@@
Although $M$ could relate to $S(t)$ and $\boldsymbol{X}$, we do not include them in the $M$ for simplicity of notation. Let $S^{\prime}(1-t)$ denote M(S(1t))M(S(1-t)) and $p_{\epsilon}$ denote the pdf of $\epsilon$. Throughout, for simplicity of notation we let $p\big(Y^{\prime}(t)=y)$ denote the pdf of $Y^{\prime}(t)$ evaluated at $y$, with similar notation for $S(t)$. The the conditional distribution of $Y^{\prime}(t)$ given $S(t)$ and $\boldsymbol{X}$ is
$\begin{split}p\big(Y^{\prime}(t)=y\mid\boldsymbol{X},S(t)\big)&=\intop\nolimits p\big(Y^{\prime}(t)=y\mid S(1-t)=s_{1-t},S(t),\boldsymbol{X}\big)dP_{S(1-t)\mid S(t),\boldsymbol{X}}\big(s_{1-t}\big)\\ &=\intop\nolimits p_{\epsilon}\big(y-f_{t}(S(t),M^{-1}(S(1-t)),\boldsymbol{X})\big)dP_{S(1-t)\mid S(t),\boldsymbol{X}}\big(s_{1-t}\big)\\ &=\intop\nolimits p_{\epsilon}\big(y-f_{t}(S(t),M^{-1}(s^{\prime}_{1-t}),\boldsymbol{X})\big)dP_{S^{\prime}(1-t)\mid S(t),\boldsymbol{X}}\big(s^{\prime}_{1-t}\big)\\ &=\intop\nolimits p_{\epsilon}\big(y-f_{t}(S(t),s_{1-t},\boldsymbol{X})\big)dP_{S(1-t)\mid S(t),\boldsymbol{X}}\big(s_{1-t}\big)\\ &=p\big(Y(t)=y\mid\boldsymbol{X},S(t)\big).\end{split}$
@@
A.2.2 Proof of an equivalent expression for Assumption 8